A uniform, solid sphere of radius 4.50 cm and mass 4.50 kg starts with a purely translational speed of 4.00 m/s at the top of an inclined plane. The surface of the incline is 1.50 m long, and is tilted at an angle of 21.0∘ with respect to the horizontal. Assuming the sphere rolls without slipping down the incline, calculate the sphere's final translational speed ?2 at the bottom of the ramp.
Answer #1

Change in kinetic energy will be given by:

dPE = dKE = KEf - KEi

KEi = kinetic energy at top = 0.5*m*v1^2

KEf = KE_trans + KE_rot = (1/2)*m*v2^2 + (1/2)*I*w^2

I = moment of inertia of solid sphere = (2/5)*M*R^2

w = angular speed = v2/R

R = radius of sphere = 4.50 cm = 0.045 m

m = mass of sphere = 4.50 kg

KEf = (1/2)*m*v2^2 + (1/2)*(2/5)*(m*R^2)*(v2/R)^2

KEf = (1/2)*m*v2^2 + (2/5)*m*v2^2 = (0.70)*m*v2^2

dPE = change in potential energy = m*g*h

h = initial height of solid sphere = 1.50*sin(21.0 deg)

v1 = initial tangetial speed = 4.00 m/s

So,

4.50*9.81*1.50*sin(21.0 deg) = 0.70*4.50*v2^2 - 0.5*4.50*4.00^2

4.50*9.81*1.50*sin(21.0 deg) + 0.5*4.50*4.00^2 = 0.70*4.50*v2^2

4.50*9.81*1.50*sin(21.0 deg) + 0.5*4.50*4.00^2 = 0.70*4.50*v2^2

v2 = sqrt[(4.50*9.81*1.50*sin(21.0 deg) + 0.5*4.50*4.00^2)/(0.70*4.50)]

v2 = 4.35 m/sec. = sphere's final translational speed at the bottom of the ramp